OpenStudy (anonymous):

Find the value of x in this proportion. 9:x = x:4

OpenStudy (lgbasallote):

change to fractions \[\frac 9x = \frac x4\] cross multiply

OpenStudy (anonymous):

x=36?

OpenStudy (lgbasallote):

no...x^2 = 36

OpenStudy (anonymous):

it would be x^2 =36 multiply by 4x to both sides

OpenStudy (lgbasallote):

x times x is x^2 remember?

OpenStudy (lgbasallote):

now solve for x by taking the square root \[x^2 = 36\]

OpenStudy (anonymous):

x=6

OpenStudy (lgbasallote):

correct

OpenStudy (lgbasallote):

or could be -6 too

OpenStudy (lgbasallote):

\[\sqrt{36} = \pm 6\]

OpenStudy (anonymous):

oohhh i get it now thank u

OpenStudy (lgbasallote):

welcome ^_^